### v1

# Reinterpreting Landauer conductance, solving the quantum measurement problem, grand unification

Abstract
In a series of recent papers we have proved rigorously that time travel is a reality and very much feasible by using quantum mechanical processes. There are plenty of indirect experimental support untill a direct experiment is conducted. The process crucially depend on the reality of a local time as well as a local partial density of states (LPDOS) that can become negative very easily in the quantum regime of mesoscopic systems. Mesoscopic systems are small enough to allow us to experimentally access the intemediate regime between the classical and quantum worlds. This LPDOS is in every sense a hidden variable in quantum mechanics that does not show up in the axiomatic framework of quantum mechanics. It can be infered through physical clocks obeying quantum dynamics and can be rigorously justified from the properties of the Hilbert space that is uniquely isomorphic to the complex plane. Therefore one can naturally guess that LPDOS will have something important to say about quantum measurement as well as the unification of classical and quantum laws. We therefore undertake the exercise to show that LPDOS can very much allow us to re-interpret the enormously successful phenomenological Landauer-Buttiker formalism for mesoscopic systems and put it on firm theoretical ground as a bridge between classical and quantum mechanics, thereby unifying them. Essentialy the local time calculated quantum mechanically can dialate exactly like the proper time of relativity and be consistent with the coordinate time of relativity. Also the measured conductance of mesoscopic samples is a deterministic quantum measurement outcome from a linear superposition of states, essentially because of LPDOS, which solves the quantum measurement problem. For this we analyze the three probe conductance formula in details and give our arguements for the general case.
1 Introduction
Mesoscopic regime is that in between the quantum and classical regime due to the small size of the samples and the low temperatures at which these systems are subjected to. Recent advances in fabrication techniques has allowed us to make high mobility two dimensional electron gas (2DEG) on which lithographic techniques can be applied to fabricate a small quantum system connected to leads that connect the quantum system to the classical world [1] and that is an essential part of these systems. These leads are labelled $\alpha$, $\beta$, $\gamma\;\;\cdots$, etc. One can make the systems so small that at low temperatures the de Broglie wavelength of the electrons become comparable to the sample size and we can look into quantum effects that are not subject to thermalization and ensamble averaging. Thus we get, what we call a general mesoscopic system that help us probe deep into the interplay of classical and quantum laws. We call it general because such a mesoscopic system becomes a true ambassador to understand the intermediate regime that lies between the quantum and classical laws. Starting from such a system, we can find variants of it to probe other problems in physics that may tend towards the classical or quantum regimes.
Motivated by the developments in fabrication of such samples, there was a flury of theoretical activity, that culminated in the celebrated Landauer-Buttiker formalism to understand these small systems. The Landauer conductance formula for example, is extremely successful in giving the conductance of such small systems, that could even be a molecule connected to leads. However, conceptual problems remain and we describe below a re-interpretation of the Landauer conductance, from the view point of our recent works [2] and mentioned in the next paragraphs.
[…]
With such a local state one can obviously ask the question if it has anything to do with quantum measurements. Buttiker et al tried to obtain a direct relation between conductance (a measurable quantity) formulas for mesoscopic systems and local partial density of states (LPDOS). They could successfully do it in the semi-classical regime but it did not work for quantum regimes where $\rho_{lpd}$ can become negative. So the notion of a local-state and local time in the quantum regime remained unsolved. A local time, if it exists in quantum regime, may shed light on how general relativity and quantum mechanics can be reconcilled. These are the two issues, that we will probe in this work. Namely, how the local time reconciles quantum mechanics and relativity and how the local state help solve the quantum measurement problem. The motivation is due to the fact that we could prove [2] that negativity of $\rho_{lpd}$ is physical and corresponds to a wave-packet travelling back in time. Since a wave-packet can also carry information or signal, this raises some philosophical issues. Hence a more down to earth experimental verification will be of great help and that is also what will be given in this work.
We will first re-interpret Landauer-Buttiker conductance formulas to show that it works because of the reality of local state and a local time in a quantum system. We will then give examples to show that what ever works in semi-classical regimes, work identically in the quantum regime. We will also show why Buttiker et al did not get it in the examples they considered.
2 Re-interpreting Landauer-Buttiker formalism
Landauers approach through DOS vs our approach of using freedom of normalization:
Figure 1: Fig. 1. A 1D quantum scatterer is sandwiched between two classical reservoirs. The source reservoir is to the left and the sink reservoir is to the right of the two wavy lines in the figure. The scatterer in between is shown in the form of a finite square barrier, whose transmission amplitude is discussed in almost all text books of quantum mechanics. The horizontal axis is the 1D coordinate axis (say $x$-axis) and the vertical axis is the energy axis. The origin of the energy axis is the band bottom of the conduction band making the sample and the height of the square barrier is shown by an arrow. Such a 1D square barrier with leads can be practically made from the GaAs-AlGaAs interface by creating a narrow confinement potential in the lateral $y$ and $z$ directions. Since the reservoirs are classical, they have a statistically large number of electrons and the system in between has only quantum states that can transmit electrons from the left reservoir to the right. The two reservoirs can have different chemical potentials and different temperatures that can be seen from the Fermi distributions generally shown inside the two reservoirs.
Landauer’s original arguement starts from the diagram in Fig. 1 and asserts that it represents a finite physical system where a) the reservoirs are classical and the source reservoir injects a statistically large number of electrons into the left lead that constitute a current for which b) there is a density of states (DOS) given by $\frac{1}{hv}$. Assertion b) is troublesome and will be discussed in detail. There can be no contradiction between reality and assertion a) and allows us to use the axioms of statistical meachanics. In Ref. [2], we have used the axioms of statistical mechanics and current conservation across a junction to show that the current through the system at a given incident energy $E$ is $\frac{e_{0}}{h}|t(E)|^{2}$ where $t(E)$ is the quantum mechanical transmission amplitude for the rectangular barrier as shown in Fig. 1. In this approach we do not need to invoke the idea of a DOS to derive the conductance. Instead we use the freedom of choosing a normalization constant that conserves current accross a junction with an electrostatic potential, and this normalization constant can be interpreted as a DOS that accommodate the electrons. In Eq. 13 we give the original arguements of Landauer, who on the other hand assumed that this current is carried by electrons in quantum states for which there is a well defined DOS. By understanding the simple 1D picture in Fig. 1 we can generalize to realistic mesoscopic sytems because of the decoupling of transverse modes in the leads [1]. The rectangular barrier and 1D leads can be replaced by realistic potentials for which the calculation of transmission probability will be complicated and one can proceed with this simplest case of a rectangular barrier in mind [1].
Unanswered questions in Landauer’s phenomenological approach: Many different experiments show that the Landauer formalism and its diverse extensions to real systems with multiple leads, finite temperature, many body effects, non-linearities, etc, remain valid even for very small lead lengths, as we will discuss. Hence the assertion b) needs to be scrutinized. i) The parameter $E$ here is externaly applied incident energy and not eigen-energy defining an energy state. How the incident energy interacts with the energy states of the system has to be treated according to accepted rigorous theory. ii) One may also ask what are the well defined states for which we have a DOS? What exactly Landauer meant when he takes the DOS available to the electrons coming from the source to be $\frac{1}{hv}$. States of exactly which system? Are they the states of the isolated left lead or the states in the leads are the asymptotic states of a quantum mechanical scattering. Isolated left lead may be thought of as an infinite ballistic system although Landauer conductance works for fairly small leads and besides that contradicts the idea that the mechanics of the electrons in the system is a quantum mechanical scattering process. In formal scattering theory, one cannot talk of DOS for asymptotic states or DOS for non-asymptotic states. A quantum mechanical scattering has to consider the states of the full system rather than that of isolated parts. In quantum mechanics there is a well defined DOS for the phenomenon of scattering between asymptotic states that is given by the resolvent and Greens function where the exact scattering potential has to be taken into account. This is related to the dwell time to be discussed below (see Eq. 12), where the integration limits is not for the region of length $l$ but has to be over the entire region between the junctions with the reservoirs. In a quantum mechanical scattering, these are energy states. iii) What exactly is a state in mesoscopic system? Is it a momentum state for which we need translational invariance or an energy state for which we need absence of dissipation? Within the system, between the reservoirs, there is no dissipation snd so energy will be conserved, energy states may be taken as good quantum states, but that will contradict how Pauli exclusion principle is applied as will be discussed. iv) Besides, the term ”number of states” is sometimes misleading as one cannot count states in the continuum of states. This contradiction of terms is not a serious matter for large systems because of a mathematically rigorous prescription of regularization that can be applied to real systems but that prescription is not adaptable to finite mesoscopic systems.
Justifying the unanswered questions in Landauer’s phenomenological approach: In formal quantum mechanics, the DOS for an infinite 1D line which has a good momentum state and also a good energy state is $\frac{1}{hv}$. There are no good momentum states for neither the mesoscopic system nor the infinite 1D line with a scatterer. An infinite 1D line with a single scatterer do not have translational symmetry and hence no good momentum states. It will still have good energy states with real energies but DOS is no longer given by $\frac{1}{hv}$. Ideas we use in condensed matter system cannot be applied here because we know that in condensed matter systems an infinite periodic potential has real eigen-energies or eigen-states of the Hamiltonian operator that are extended. Also an infinite random potential has real eigen-energies that are localized. Mesoscopic phenomenon arise from finiteness of the system. Besides, in condensed matter systems, WKB approximation works well that tells us there can be a local momentum corresponding to the classical momentum. If we assume these features then we are still in the condensed matter regime and not accessing the mesoscopic regime.
Semi-classical approach: In formal quantum mechanics only infinite systems have a continuous spectrum implying that states are uncountable. Therefore we can start from some finite system to say DOS is defined as the number of states per unit energy per unit length, which is then regularized to give measure or volume of states of the full system at an energy devided by the volume of the system to mean that at a given point or cross section. Mesoscopic systems being finite by definition, do not allow this rigorous prescription of regularization and yet electrons coming from the classical reservoir has a continuously varying incident energy. The prescription followed by S. Datta [1] based on use of Dirac delta function is ad hoc. Dirac delta function is good for solving linear differential equations but it is difficult to translate what a Dirac delta function means for finite quantum systems in nature and whether it helps us bypass the standard mathematical rules of regularization by introducing a complex self-energy parameter for a finite system. For the system size tending to infinity, there is a very close connection between the so called von Neumann’s approach through Hilbert space and Dirac’s approach through bra-ket notation. One being rigorously mathematical and the other being more intuitive, with the caution that the intuitive approach may be very helpful but can overlook errors. Dirac’s approach looses its effectiveness unless space is discritized and the Hamiltonian is represented as a matrix. Thus the self-adjoint Hamiltonian operator is reduced to a Hermitian matrix to give Heisenberg’s matrix mechanics, and then the continuum limit can be imposed through a hopping parameter without following the mathematical rigour of taking limits. If one is ready to accept some parameter fitting then this method adopted by S. Datta [1] is good for numerical calculations and hopefully that is adequate in the semi-classical regime. One may state that such ad-hoc approaches cannot reconcile the theories of quantum mechanics and relativity. One may ask what role does the hopping parameter or the self energy parameter play in general theory of relativity and how do they transform. Besides as it has been shown [2] that one cannot define a momentum oprator for a general mesoscopic system in the quantum regime, canonical quantization will not work and von Neumann’s wave-mechanics is essential.
Our fix for what is a state in a mesoscopic lead for which we write a DOS? For this note the fact that given a problem, von Neumann quantum mechanics first require us to determine the Hilbert space. The Hilbert space is determined by specifying the basis-states. We can use the eigen-states of some well defined operator as the basis-states. Since there is only one Hilbert space (that is isomorphic to the complex plane), any complete orthonormal basis may work, only restrictions can come from boundary conditions or physical constraints or dimensionality of the sample, etc, that restrict the theory to a subspace.
Quantum mechanics start with the fact that nature generates a wave-packet that evolves in time, for example, the classical reservoirs inject a pulse of classical electrons into the system. Inside the mesoscopic system, the pulse has to treated as a wave-packet using time dependent Schrodinger equation (TDSE) and quantum mechanics start with this wave-packet. The wave-packet can be decomposed into partial waves that satisfy time independent Schrodinger equation (TISE) and quantum mechanics tells us how to relate the two dynamics using the probabilistic interpretation and uncertainty principle. Essentially many events of pulse propagation averages to give the scattering probabilities and all other physical effects that one can get from the steady beam scattering. A steady beam or a stationary state can be only viewed as a limiting situation of uncertainty in momentum going to zero and taking this limit has to be treated with extreme mathematical rigorousness. This cannot be done by assuming that a Kronecker delta smears into a Dirac delta function as is often done [1] (see page 97 therein).
Finally, a wave-packet (also a Hilbert space element at an instant of time) is what propagates and has a time dependent equation of motion. The wavelets of the wave-packet have wave-vectors that vary continuously. While discrete states can be counted, continuum of states has a density. We can completely ignore the notion of DOS to arrive at local time recorded by a physical clock (a Larmor clock) [3]. Larmor clock devided by $\hbar$ give DOS which is completely consistent with the DOS we define quantum mechanically [3]. This DOS can be seen as how many wave-vector values are needed to make a full topological circle in a complex plane, say of the complex transmission amplitude $t$ ($Im(t)$ versus $Re(t)$) [2], as will be explained below. Thus the idea of a quantum number is replaced by winding number in this complex plane. Thus there is something like a local quantum number at the point r which is more appropriately a winding number $\delta\theta_{r}$ or $\delta\theta_{t}$ for which there is a local DOS (called local partial density of states or LPDOS)
$\rho_{lpd}^{r}=-\frac{1}{2\pi e_{0}}\frac{\delta\theta_{r}}{\delta U(\textbf{r})}\;\;\mbox{and}\;\;\rho_{lpd}^{t}=-\frac{1}{2\pi e_{0}}\frac{\delta\theta_{t}}{\delta U(\textbf{r})}$ (1)
for the reflected and the transmitted electrons, respectively. Here $\theta_{r}$ and $\theta_{t}$ are the phases of the complex reflection amplitude $r$ and transmission amplitude $t$ , respectively, that change with $k$ and hence $U(\textbf{r})$, $U(\textbf{r})$ being the electro-static potential at the point r and the derivatives are functional derivatives. We have shown that this can be justified through physical clocks as well as functional analysis, without relying on operators in quantum mechanics and can be seen in more detail in [2]. At low energy $U(\textbf{r})$ is classical with no quantum corrections while only the particles are fully subject to quantum mechanics. However, in appropriate conditions, $U(\textbf{r})$ can exhibit quantum fluctuations for which there is a well established theory. This local DOS in Eq. 1 is specifically called local partial density of states (LPDOS) as it holds for only reflected electrons or only transmitted electrons. The objective reality of LPDOS can be established from functional analysis without demanding quantum states. It is consistent with quantum mechanics, but it cannot be found within the axiomatic framework of quantum mechanics [2]. The winding numbers gradually evolve into discrete quantum numbers of appropriate operators, as we pinch off the system and isolate it from external incidence from a classical reservoir.
In quantum mechanics ”states” is an ill defined concept and science deals with such ill defined concepts in terms of properties or qualities for which we use an appropriate adjective to get a qualified state. One can cite other examples like we do not understand what is time but we can define diffusion time, or first passage time etc. We do not understand causality but we can define space-time-causality, or interventionalist’s-causality. Our local partial density of states (LPDOS) defined in Eq. 1, are just states and not to be hyphenated with any operator but consistent with appropriate states of a quantum operator, depending on the problem at hand. This is further elaborated as we proceed.
Pauli exclusion Inside the mesoscopic system, the Landauer-Buttiker formalism assumes Pauli exclusion principle for each wave-vector separately for transmitted and reflected electrons that does not necessarily define a momentum-state or an energy-state and we can justify this assumption as done below from LPDOS. Let us consider the reflected electrons first. We understand that the LPDOS also define a time spent [2] by the reflected particle at the coordinate r that is given by
$\Delta\tau_{lpt}^{r}=-\frac{\hbar}{e_{0}}\frac{\delta\theta_{r}}{\delta U(\textbf{r})}$ (2)
This time is consistent with the mesoscopic Larmor clock which in turn is consistent with Lorentz’s transformation as
$U\rightarrow\gamma U\;\;\mbox{and}\;\;\gamma=\frac{1}{\sqrt{1-v^{2}/c^{2}}}$ (3)
and gives proper time (Larmor clock is body frame clock) correctly that is less than the coordinate time difference between two classical points in the classical reservoirs. Every point inside the sample can thus be assigned a coordinate time (obtained from Eqs. 2 and 3 with $v\rightarrow 0$) that is in the laboratory frame and from this frame the body clock runs slow for $v\neq 0$. In analogy, a typical Stern-Gerlach set up where spins precess, every point inside the system can have three spatial coordinates due to three perpendicular spatial axes (the classical magnetic field can be oriented in three perpendicular directions) while for each independent spatial direction, the two spin degrees of freedom define an internal orthogonality in the body frame. Following Eqs. 2 and 3, one can assign three spatial coordinates and one time coordinate to every point in the sample and there is an additional internal body frame clock that records a proper time that is consistent with relativity. This is also true for the classical Larmor clock for which one can see [4]. Hence the local state at the point r is consistent with quantum mechanics as well as Lorentz’s transformation and therefore by the statement of spin statistics theorem, we can apply Pauli exclusion principle to the local state defined by a winding number. The PDOS for an extended region can be found by integrating the LPDOS over this region and the local time adds up to give a propagation time over an extended region. The same can be said for the transmitted modes.
$\Delta\tau_{lpt}^{t}=-\frac{\hbar}{e_{0}}\frac{\delta\theta_{t}}{\delta U(\textbf{r})}$ (4)
To find the LPDOS at a point r in the right lead (say) we have to note that $\theta_{t}$ is the same at all points on the right lead and the phase shift to be considered at a distance of $l$ from the scatterer is
$\theta_{t}+kl$ (5)
were $l$ is the length between the end point of the square barrier potential and the point r of the right lead. Electrostatic potential $U$ in the leads is given by the bottom of the conduction band and is constant. Therefore, in the leads,
$k=\sqrt{\frac{2m_{0}}{\hbar^{2}}(E-U)}$ (6)
Comparing Eqs. 1 and 6, we can say that to talk about the available states in the right lead for the transmitted electrons, we have to apply an infinitesimal $U(\textbf{r})$ at the point r on the right lead (see Eq. 1) and then integrate over r in the right lead. Hence, in the right lead, from Eq. 6, one gets
$\frac{d}{dE}\equiv-\frac{1}{e_{0}}\underset{\mbox{right lead}}{\int}d\textbf{r}\frac{\delta}{\delta U(\textbf{r})}$ (7)
leading to the fact that electrons access a number of states in the right lead at a point and at an energy $E$ given by $\frac{1}{hv}$. This can be seen as follows.
$\rho_{pd}^{t}=\int d\textbf{r}\rho_{lpd}^{t}=-\frac{1}{2\pi e_{0}}\int d\textbf{r}\frac{\delta(\theta_{t}+kl)}{\delta U(\textbf{r})}=-\frac{1}{2\pi e_{0}}\int d\textbf{r}\frac{\delta(\theta_{t})}{\delta U(\textbf{r})}-\frac{1}{2\pi e_{0}}\int d\textbf{r}\frac{\delta(+kl)}{\delta U(\textbf{r})}=-\frac{1}{2\pi e_{0}}\int d\textbf{r}\frac{\delta(+kl)}{\delta U(\textbf{r})}$ (8)
As $\theta_{t}$ is independent of a vanishingly small $U(\textbf{r})$ in the lead acting on the partial +k+k mode, that is $\frac{\delta(\theta_{t})}{\delta U(\textbf{r})}\bigg|_{\textbf{r}\in\mbox{(right lead})}=0$. This $U(\textbf{r})$ is not meant to reflect a transmitted electron but only meant to give a derivative (like a tangent to a curve does not need to increament the curve). Note that Eq. 7 is not an equality but what it says is that increasing the incident energy for a fixed $U$ in the lead, is equivalent to decreasing the $U$ in the lead for a fixed incident energy. Therefore from Eqs. 6 and 7 and noting the fact that $k$ is constant for all r in the lead ($k$ will however have a gradiant with $U(\textbf{r})$ with $E$ constant or will have a gradient with E for $U$ constant),
$\rho_{pd}^{t}=\frac{1}{2\pi}\frac{d(+kl)}{dE}=\frac{1}{2\pi}\frac{d(+kl)}{d(+k)}\frac{d(+k)}{dE_{(+k)}}=\frac{1}{2\pi}\frac{d(+kl)}{d(+k)}\frac{d(k)}{dE_{(k)}}=\frac{1}{2\pi}\frac{d(+kl)}{d(+k)}\frac{1}{\frac{dE}{dk}}=\frac{1}{2\pi}\frac{l}{\frac{d}{dk}\frac{\hbar^{2}k^{2}}{2m_{0}}}$
$=\frac{2m_{0}l}{2\pi\hbar^{2}}\frac{1}{2k}=\frac{m_{0}l}{2\pi\hbar^{2}k}=\frac{l}{h}\frac{1}{\frac{\hbar k}{m_{0}}}=\frac{l}{hv}$ (9)
The factor $\frac{1}{2}$ comes because when we write the dispersion relation as $E=\frac{\hbar^{2}k^{2}}{2m}$ then both $\pm k$ states are taken into account. The transmitted electrons are in the +k+k partial mode and therefore,
$\rho_{pd}^{t}=\frac{l}{hv}$ (10)
Since LPDOS physically (not just meant) add up to give PDOS
$\rho_{lpd}^{t}=\frac{1}{hv}$ (11)
Thus from Eq. 11, we have local partial states at a point on the right lead with a density given by $\rho_{lpd}^{t}=\frac{1}{hv}$ and for the extended region of the right lead, we get partial states with a density given by $\rho_{pd}^{t}=\frac{l}{hv}$ (from Eq. 10) and both are equally physical because LPDOS is a physical object that is coordinate dependent in general although not in the leads, that add up to give PDOS. This is thus different from DOS for the states of the full system that we get from axiomatic quantum mechanics that can be mathematically devided by the volume of the system to mean something. To be noted that, $\rho_{lpd}^{t}$ is due to local time spent at a point and $\rho_{pd}^{t}$ is due to sum of local times over the region $l$, and both comply with Lorentz’s trnsformation. At every point r and we can apply Pauli exclusion principle to the mode +k+k. So the LPDOS for +k+k mode in the right lead turns out to be the same at all points on the right lead, is $\frac{1}{hv}$, and it is subject to Pauli exclusion principle. The same is similarly true for the reflected electrons with wave-vector k-k. However, for r inside the system with a constant potential barrier, $\frac{d\theta_{t}}{dE}\not\equiv\frac{1}{e_{0}}\frac{\delta}{\delta U}\theta_{t}$ because of linear superposition of $\pm k$-modes and there is no decoupling of the $\pm k$ modes due to multiple Feynman paths. For example, $\frac{1}{e_{0}}\frac{\delta}{\delta U(\textbf{r})}\theta_{t}$ will be r dependent, depending on how close we are to the edge of the constant potential barrier. In the leads, this superposition is destroyed by the classical reservoirs [2]. For a non-rectangular barrier, say for a wavy potential barrier, inside the system there is no idnetifiable $\pm k$ states, $\frac{d\theta_{t}}{dE}\not\equiv\frac{1}{e_{0}}\frac{\delta}{\delta U}\theta_{t}$, but the leads can again be treated exactly like that of the rectangular barrier.
Eqs. 8, 9, 10 and 11 are consistent with a dwell time in a section of length $l$ in the leads that can be defined within the axiomatic framework of quantum mechanics. For the electrons incident from the left reservoir into the left lead we can write the wavefunctions as $\psi(k,x)=e^{ikx}$ and hence the dwell time as
$\frac{1}{v}\int_{0}^{l}dx|\psi(k,x)|^{2}=\frac{l}{v}$ (12)
and devided by $h$ a DOS. Here $l$ can be aribitrarily small or large. The dwell time can be found within the axiomatic framework of quantum mechanics as in a general situation (including the situation in a lead where superposition of states is destroyed by a reservoir) $\psi(k,x)$ inside the system does not have any information of from where the electrons come and where they go, but not LPDOS.
Outline: While the general view is that quantum mechanical laws are the fundamental laws of nature and classicality emerges, we adopt the view that a quantum event can only be defined between two classical events in the reservoirs and the quantum laws can coexist with classical laws without any contradiction. This has some obvious but significant consequences and solves some long standing problems in physics. First is that theory of relativity and quantum mechanics can co-exist consistently and there is no immediate need to quantize gravity, besides the fact that $U(\textbf{r})$ has well known quantum behavior. That is because now in quantum mechanics, like in theory of relativity, we can relate time intervals to signal propagation time. It was shown in ref. [2] that this signal is carried by a propagating localized undispersed quantum mechanical wave packet. We can clearly obtain a coordinate time and a proper time within the framework of quantum mechanics. The points inside the classical reservoirs have local clocks giving their respective local times or coordinate time that is consistent with mass distribution in synchronizing the classical local clocks. A signal can propagate from one reservoir to another reservoir according to the laws of quantum mechanics and the time delay defines a proper time that can dialate and contract and can be different from local times clocked by the classical clocks inside the reservoirs. The quantum system gets a space-time formulation in the sense that every point inside the quantum system has three spatial coordinates and a time coordinate. This time coordinate or coordinate time ($v\rightarrow 0$ in Eq. 3) is consistent with proper time that will be recorded by the body clock of a particle or wave-packet traversing the points and that in turn defines an un-hypheneated local quantum state that does not need a Hilbert space. Thus we get grand unification of quantum and classical laws at low energies rather than at high energies.
One can present three very strong evidences in favor of this view. First is that this is supported by the enormously successful Landauer-Buttiker formalism as will be shown in sub-section 2.1. Second is that it leads to a natural understanding of what happens in a measurement on a quantum system, that is, solves the measurement problem in quantum mechanics as will be shown in sub-section 2.2. Third is that it leads to a definition of time in a quantum system that is fully consistent with STR and GTR, and this time does not crop up in the von Neumann’s axiomatic framework of quantum mechanics and has been missed so far. These points will be exemplified through calculations below. This also implies that unlike in von Neumann quantum mechanics or any other version, we do not have a quantum dynamics extending from -$\infty$ to $\infty$ and that can naturally remove certain divergences.
2.1 Application to Landauer-Buttiker conductance formulas
The above mentioned idea of exclusion principle applied to local partial states will be now used to re-describe Landauer-Buttiker formalism. We will apply exclusion principle to local partial states at zero temperature as well as to finite temperatures. It will be thus clear that Fermi-Dirac statistics applies to local partial $k$ modes and that finds strong experimental support in the overwhelmingly successful Landauer-Buttiker conductance formulas and is explained in detail below. There are no states and no Fermi-Dirac distribution function in the reservoirs unlike what is presented in Fig. 1 as a standard practice. One typically draws a Fermi-Dirac distribution function in the reservoirs as shown in Fig. 1, and as can be seen in the book by S. Datta [1]. This is in fact un-necessary or may be a matter of convinience, but could also be related to a typical mindset that at very low temperatures the reservoirs become quantum mechanical. The general belief is that quantum mechanics is fundamental and classicality emerges. As a consequnce of this general belief, one may think that at zero temperature the whole universe will become quantum mechanical with the caveat or circular logic that zero temperature can never be reached. In this general belief system one draws a Fermi-Dirac distribution inside a reservoir. Mesoscopic experiments are done at extremely low but finite temperatures such that $k_{B}T$ is less than sub-band spacing in the leads and one should never consider the reservoirs to be at zero temperature. A classical reservoir can have a chemical potential and temperature which are just parameters and not related to a Fermi-Dirac distribution of electrons. Fermi-Dirac distribution can be however, considered once the electrons enter the leads and the system that are quantum mechanical. In the region between the reservoirs we have a scattering problem that is described in general in the framework of von-Neumann quantum mechanics [2] (see ch 4 therein) that can be easily adapted for the 1D case of Fig. 1. In the Landauer’s approach, the mathematics of quantum mechanics connects to currents in nature through $ne_{0}v$ where $n$ is the number (countable) of carriers which in this case is electrons, $e_{0}$ is the charge of an electron and $v$ is its velocity.
Two probe: The left reservoir injects a current into the mesoscopic sample for which we have discussed that the available states at any point for +k+k (see Eq. 11) is $\rho^{t}_{lpd}=\frac{1}{hv}$, where $v=\frac{\hbar k}{m_{0}}$, $m_{0}$ being the mass of an electron. For right moving electrons $k$ is positive, while negative for left moving electrons. So total injected current in the left lead, at a point close to the reservoir, between $\mu_{1}$ and $\mu_{2}$, that is in the energy range $dE=(\mu_{1}-\mu_{2})$ will be
$dJ_{in}=(dn)e_{0}v=\frac{dn}{dE}e_{0}v(dE)=\frac{1}{hv}e_{0}v(dE)=\frac{e_{0}}{h}(\mu_{1}-\mu_{2})$ (13)
Therefore the differential current that pass through the sample to reach the sink reservoir is
$dJ=\frac{e_{0}}{h}|t(E)|^{2}(\mu_{1}-\mu_{2})$ (14)
If there is a current between a source and a sink then one can easily define a conductance as will be done below. In a typical two probe conductance measurement wherein the voltage difference is measured between the two classical reservoirs, the conductance will be
$G_{two-probe}^{zero-temperature}=\frac{\frac{e_{0}}{h}|t(E)|^{2}(\mu_{1}-\mu_{2})}{\frac{\mu_{1}-\mu_{2}}{e_{0}}}=\frac{e_{0}^{2}}{h}|t(E)|^{2}$ (15)
If we consider spin degeneracy for positive $k$ then DOS will be $\frac{2}{hv}$ and this formula becomes
$G_{two-probe}^{zero-temperature}=\frac{\frac{2e_{0}}{h}|t(E)|^{2}(\mu_{1}-\mu_{2})}{\frac{\mu_{1}-\mu_{2}}{e_{0}}}=\frac{2e_{0}^{2}}{h}|t(E)|^{2}$ (16)
Thus we have adopted the fact that inside the system, when $k_{B}T$ is less than sub-band spacing due to lateral confinement, we can apply zero temperature quantum mechanics. At zero temparature, all states in the conduction band is filled below $\mu_{2}$ and are also equilibrated which can also be seen as the principle of detailed balance as will be explained below. The conduction band is only in the lead because of the way in which the samples are fabricated from GaAs/AlGaAs interface. The reservoir has no Fermi energy but even classical reservoirs have a chemical potential and temperature. When the electron enters the lead then it gets a chemical potential which is also the Fermi energy.
We will now proceed to account for finite temperature effects, voltage differences and role of multiple leads. All these formulas has been rigorously shown to agree with experiments [see [1]]] inspite of the severe conceptual problems discussed above. We remove these conceptual problems and thereby show that this naturally solves other problems in physics.
Finite tempreature : Now let us see what may happen at finite temperature when states above $\mu_{1}$ and $\mu_{2}$ has a finite probability to be occupied. The left lead will be in equilibrium with the left reservoir. Say there are $\frac{dn}{dE}$ number of relevant modes at any (all) point in the left lead and not all of them are occupied. Number of electrons at any point in the left lead will be given by the number of occupied modes, that is $\frac{dn}{dE}f(E,\mu_{1})$ where $f(E,\mu_{1})$ is the Fermi-Dirac distribution function. Similarly, number of electrons in right lead is $\frac{dn}{dE}f(E,\mu_{2})$.
So number of candidate electrons on the left of the scatterer that wants to cross to the right is
$\frac{dn}{dE}\bigg(f(E,\mu_{1})-f(E,\mu_{2})\bigg)$ (17)
Occupied +k+k modes in left lead contribute positively, while occupied +k+k modes in right lead are subtracted to get Eq. 17. For the +k+k modes only, the relevant DOS is $\frac{dn}{dE}=\frac{1}{hv}$ in both leads. Number of them that actually cross to the right is
$\frac{dn}{dE}\bigg(f(E,\mu_{1})-f(E,\mu_{2})\bigg)|t(E)|^{2}$
Although we call it a number, it seems to have a dimension of $\frac{1}{E}$ as $df$ and $|t|^{2}$ are numbers. This gets rectified at the next step where we write current that has a dimension of $\frac{e_{0}}{h}$. Current $ne_{0}v$ through the system will therefore be
$\frac{dn}{dE}\bigg(f(E,\mu_{1})-f(E,\mu_{2})\bigg)|t(E)|^{2}e_{0}v$ (18)
$\frac{1}{hv}\bigg(f(E,\mu_{1})-f(E,\mu_{2})\bigg)|t(E)|^{2}e_{0}v$
$\frac{e_{0}}{h}\bigg(f(E,\mu_{1})-f(E,\mu_{2})\bigg)|t(E)|^{2}$
$\frac{e_{0}}{h}\frac{\partial f}{\partial\mu}d\mu|t(E)|^{2}=\frac{e_{0}}{h}(-\frac{\partial f}{\partial E})(\mu_{1}-\mu_{2})|t(E)|^{2}\;\;\mbox{as}\;\;\frac{\partial f}{\partial\mu}=-\frac{\partial f}{\partial E}$
Total current at finite temperature will be
$I_{two-probe}^{finite-temperature}=\int_{0}^{\infty}\frac{e_{0}}{h}(-\frac{\partial f}{\partial E})|t(E)|^{2}(\mu_{1}-\mu_{2})dE$ (19)
$I_{two-probe}^{finite-temperature}=\frac{e_{0}}{h}(\mu_{1}-\mu_{2})\int_{0}^{\infty}dE(-\frac{\partial f}{\partial E})|t(E)|^{2}$
For the two probe conductance one has to measure the voltage difference between the two classical reservoirs, which again turns out to be
$G_{two-probe}^{finite-temperature}=\frac{I_{two-probe}^{finite-temperature}}{\frac{\mu_{1}-\mu_{2}}{e_{0}}}=\frac{e_{0}^{2}}{h}\int_{0}^{\infty}dE(-\frac{\partial f}{\partial E})|t(E)|^{2}$
Let us take the zero temperature limit wherein $f(E)\rightarrow\theta(E_{f}-E)$ and $(-\frac{\partial f}{\partial E})\rightarrow\delta(E-E_{f})$ to get
$G_{two-probe}^{zero-temperature}=\frac{e_{0}^{2}}{h}|t(E_{f})|^{2}$ (20)
Since the Fermi energy or the incident energy is determined by the left reservoir we may replace $E_{f}$ by $E$ to get
$G_{two-probe}^{zero-temperature}=\frac{e_{0}^{2}}{h}|t(E)|^{2}$ (21)
which is same as Eq. 15.
Assuming an exact LPDOS of $\frac{1}{hv}$ in the finite leads play a further role when we discuss charge neutrality in the leads as will be presented below. It assumes that there are wave-vectors +k+k and k-k to which we can apply Pauli exclusion principle separately. The $\pm k$ wave-vectors naturally come from the wavelets making the wave-packet which is the central object of study in quantum mechanics. But assuming that k-k states are occupied only if there is reflection (as will be done below) from the scatterer etc, and apply Pauli exclusion principle for the $\pm k$ modes separately without questioning whether they define quantum states is ad hoc. While developing a conductance formula it is not convinient to talk of energy states, and although one may argue that energy is conserved between the reservoirs, $\pm k$ belong to the same quantum state (energy state), which would mean that if +k+k state is occupied then k-k state cannot be occupied. In fact that the current through a system is given by the transmission probability through the system was known much before Landauer and it was thought, that happens only when we have sparsely filled band or very low transmission probability such that multiple electrons competing for a state is negligible. One may see [5] or can also see the more recent work of [6]. Surprise came with the success of Landauer approach for densely filled bands, transmission probability approaching the maximum limit etc. A lot of space has been given in S. Datta’s book to discuss this matter but a confusion can only be expressed in a confusing manner as can be seen on page 93 of [1].
Towards a four probe Landauer conductance formula at zero temperature: Note that the central object in quantum mechanics that is a wave-packet can be decomposed into wavelets and that naturally give four different wavelets. They are +k+k mode moving forward in time, +k+k mode moving backward in time, k-k mode moving forward in time and k-k mode moving backward in time. In this section we are only talking about $\pm k$ modes in a wave-packet moving forward in time. A +k+k mode moving in forward time is physically indistinguishable from a k-k mode moving in backward time and to keep the arguements simple we only consider forward moving modes. If backward moving modes manifests (for example it does in the viscinity of a Fano resonance when the entire wave-packet can move backwards in time [2]) the arguements can be accordingly adjusted easily. Once again by zero temperature we only mean $k_{B}T$ is less than the subband spacing in the leads (see ch4 in [2]). For the two probe conductance, the voltage difference betwee the two classical reservoirs are considered. For a four probe conductance we need the voltage difference between the two leads and the leads are quantum in nature, thus leading to complicacies related to measurements in quantum mechanics. Theoretically we can always calculate a voltage difference between the two quantum leads based on the countability of electrons, and so we proceed to calculate the chemical potentials $\mu_{A}$ of left lead and $\mu_{B}$ of right lead, at zero temperature. For a completely opaque barrier we will get the situation $\mu_{A}=\mu_{1}$ and $\mu_{B}=\mu_{2}$. For a barrier that transmits, we will get a situation $\mu_{1}>\mu_{A}>\mu_{B}>\mu_{2}$ as shown in Fig. 1. First consider the right lead where we only have to consider the +k+k modes for which the number of available states at a point is given by the LPDOS, and for the relevant current carrying energy window the number of electrons is $\frac{dn}{dE}(\mu_{1}-\mu_{B})$. Of all these available states, a state will be occupied only if it receives an electron transmitted across the scatterer from the left reservoir. Therefore the number of occupied states in right lead, between $\mu_{B}$ and $\mu_{1}$ is
$n^{oc}_{rl}=\frac{dn}{dE}(\mu_{1}-\mu_{B})|t(E)|^{2}$ (22)
There will be also some unoccupied states between $\mu_{2}$ and $\mu_{B}$ that can be found as follows. The total number of available states in this range is
$2\frac{dn}{dE}(\mu_{B}-\mu_{2})$
The factor 2 comes because in this range of energy we have to consider right moving +k+k states as well as left moving k-k states. If we imagine a situation when all these states are unoccupied and then some of the +k+k states will be filled up by electrons transmitted across the scatterer from the left. Which means, of the $\frac{dn}{dE}(\mu_{B}-\mu_{2})$ right moving states in the right lead, the number of unoccupied ones will be
$\frac{dn}{dE}(\mu_{B}-\mu_{2})-\frac{dn}{dE}(\mu_{B}-\mu_{2})|t|^{2}$
All the k-k states will be unoccupied because the only way that these states can be occupied is if they recieve an electron from the right reservoir and at zero temperature this cannot happen in the energy interval $\mu_{B}-\mu_{2}$. Any transmitted electron from left or reflected electron from the right can only go to a +k+k state. Which means all the $\frac{dn}{dE}(\mu_{B}-\mu_{2})$ left moving states in this energy range will be unoccupied in the right lead. So the total number of unoccupied states in right lead will be
$n^{un-oc}_{rl}=2\frac{dn}{dE}(\mu_{B}-\mu_{2})-\frac{dn}{dE}(\mu_{B}-\mu_{2})|t(E)|^{2}$ (23)
Let us now consider the left lead. Number of occupied local states at a point between $\mu_{A}$ and $\mu_{1}$
$n^{oc}_{ll}=\frac{dn}{dE}(\mu_{1}-\mu_{A})+\frac{dn}{dE}(\mu_{1}-\mu_{A})|r(E)|^{2}=\frac{dn}{dE}(\mu_{1}-\mu_{A})(1+|r(E)|^{2})$ (24)
The first term is for +k+k modes and all of them are occupied by electrons coming from left reservoir. The second term is for k-k modes that are occupied only if there is reflection whose probability is $|r(E)|^{2}$. Number of unoccupied stsates in left lead between $\mu_{2}$ and $\mu_{A}$ will be
$n^{un-oc}_{ll}=2\frac{dn}{dE}(\mu_{A}-\mu_{2})-\frac{dn}{dE}(\mu_{A}-\mu_{2})(1+|r(E)|^{2})$ (25)
The first term gives the total number of local states at a point. From them if we subtract the occupied ones then we will get the number of unoccupied states. So the second term can be understood by looking at Eq. 24. Charge neutrality requires
$n^{oc}_{ll}=n^{un-oc}_{ll}\;\;\mbox{and}\;\;n^{oc}_{rl}=n^{un-oc}_{rl}$ (26)
Since we are equating occupied states to unoccupied states to get charge neutrality (an occupied state having arbitrary number of electrons and unoccupied state with no electrons cannot give charge neutrality), again we are considering that a state can be occupied by one electron only. Solving Eq. 26 one can get
$\mu_{A}-\mu_{B}=(\mu_{1}-\mu_{2})|r(E)|^{2}$ (27)
If this voltage difference can be measured then we will get a conductance that can be obtained from Eqs. 14 and 27
$G^{hypothetical}_{four-probe}=\frac{\frac{e_{0}}{h}|t(E)|^{2}(\mu_{1}-\mu_{2})}{\frac{\mu_{A}-\mu_{B}}{e_{0}}}=\frac{e_{0}^{2}}{h}\frac{|t(E)|^{2}}{|r(E)|^{2}}$ (28)
It is labeled hypothetical because this conductance is purely theoretical and what will be measured cannot bypass the measurement problem of quantum mechanics. That is analogous to eigenstates of an atom that can be calculated but cannot be observed because any attempt to observe will disturb the state. This also tells us why a conductivity of a mesoscopic sample can be calculated from Eq. 28 but cannot be observed. This also makes it clear that the Hamiltonian of a system cannot tell us the conductance. As to what is actually observed or measured in several experiments on such systems, we provide the following analysis given by Buttiker.
But before we go to Buttikers analysis, it is worth mentioning that Eq. 28 does tell us something very important. From Eq. 28 we can write for the resistance
$R^{hypothetical}_{four-probe}=\frac{h}{e_{0}^{2}}\frac{1-|t(E)|^{2}}{|t(E)|^{2}}$
$R^{hypothetical}_{four-probe}=\frac{h}{e_{0}^{2}}\frac{1}{|t(E)|^{2}}-\frac{h}{e_{0}^{2}}$
Now we can see from Eq. 15 that for a perfect conductor of length $l$, there is no reflection and so $|t(E)|^{2}=1$, implying there is a contact resistance of $\frac{h}{e_{0}^{2}}$ where the classical source reservoir meets the lead. So to this hypothetical sample resistance if we add this contact resistance then we get back Eq. 15.
$R^{hypothetical}_{four-probe}+\mbox{real contact resistance}=\frac{h}{e_{0}^{2}}\frac{1}{|t(E)|^{2}}-\frac{h}{e_{0}^{2}}+\frac{h}{e_{0}^{2}}$
[…]
The contact resistance and the sample resistance are in series and just add up like resistors in series. The contact resistance can always be measured by measuring the resistance of a perfect conductor but the sample resistance $R^{hypothetical}_{four-probe}$ cannot be measured separately as we will show below, but can be only thus infered from Eqs. 15 and 28.
Measuring the two probe resistance and the contact resistance always involves two probes with two contacts. Thus we need at least two contacts to talk of a resistance, and hence there is no way to know if the contact resistance is equally shared by the two contacts or it appears only at the source contact or only at the drain contact. Ref. [7] argue that the drain contact is reflectionless and hence resistance less. Their idea is that drain contact resistance can be thought of as a current carrying narrow wire with well spaced levels injecting the current into a wide wire with close by levels. This suggests the idea that when the close by countable levels come very close then they form a continuum of levels. It is mathematically wrong to say that countable states can come close together to form the continuum. Besides, continuum of states do not necessarily make classicality emerge and it is an unsolved problem in physics as to how classicality emerges in the reservoirs. It will be theoretically impossible to show or argue that the wide conductor can gradually become a classical reservoir. An experiment can only measure the two probe conductance of a sample that consist of a narrow conductor and a wide conductor joined together between two contacts, even though the narrow-wide junction could be fully transmitting. These scientific explorations come from the idea that the universe is fundamentally quantum in nature and classicality emerges. For a comprehensive understanding of the experimentally proven Landauer-Buttiker formalism we have to discard this view.
Noting the fact that the contact resistance is in terms of fundamental constants that can be measured to an accuracy of one part in a billion we can not assume mesoscopic phenomenon is an approximation to condensed matter phenomenon. In the next sub-section, we will be specifically discussing a regime in which WKB approximation fails miserably and we will show that in that regime, there is a new origin to local states compatible with Lorentz’s symmetry and hence subject to Pauli exclusion principle. This will therefore make it obvious that classical laws can co-exist with quantum laws without contradicting each other and without having one to evolve into the other. This will also solve the quantum measurement problem. Let us see below what will be measured if we try a four probe measurement.
Multi probe conductance: We put voltage and current probes in the same footing. Which means we consider a mesoscopic sample connected to multiple leads and we have not yet fixed which lead will be used as a current probe and which as a voltage probe. They will all cause decoherence (phase randomization) in the same manner, that is in terms of causing decoherence, all leads are equivalent. Devices like voltmeter or ameter are classical devices, reading off the system after the necessary decoherence has happened. There will be no need for collapse as we will see. Collapse of wave-function will not be random and hence not a consequence of the decoherence but rather a consequence of the fact that a quantum event can only be defined between two classical events in space-time. Hence let us say that there are multiple leads attached to a mesoscopic sample that are labelled $p$, $q$, etc. Let us focus on the lead $p$ which is attached to a reservoir that has a chemical potential $\mu_{p}$. This reservoir injects at a rate of $ne_{0}v$ number of electrons into this lead at a point in the lead attached to it, that is determined by the LPDOS with Pauli exclusion principle applied to the local state at that point at an energy $E$. Thus the injected current into the mesoscopic sample is determined by the injecting lead alone, which was assumed by Landauer without any reason, and one can define an injectance of this lead with respect to the full system to which it is attached with several outgoing leads, wherein the injectance is like a resistance to this injected current [2]. Thereby, we write
$(dn)e_{0}v=\frac{dn}{dE}e_{0}vdE=\frac{1}{hv}e_{0}vdE=\frac{e_{0}}{h}dE=dJ$
Current flowing out from lead $p$ to lead $q$ will be
$J^{u}_{qp}=\frac{e_{0}}{h}|t_{qp}|^{2}(\mu_{p}-0)$
Current flowing in from lead $q$ to lead $p$ will be
$J^{u}_{pq}=\frac{e_{0}}{h}|t_{pq}|^{2}(\mu_{q}-0)$
Here the transmission from $p$ to $q$ or from $q$ to $p$ is with the other probes removed or else $J^{u}_{qp}$ and $|t_{qp}|$ will not be related. Trnasmission to other leads will contribute to $J^{u}_{qp}$. We have also seen that the two probe conductance best captures the property of the system in terms of a sample resistance that is independent of measurements along with a constant additive factor that is the contact resistance. Thus for the two probe case, when we calculate net current flowing from lead $p$ to lead $q$, this contact resistance will cancel to give us
$J_{qp}=J^{u}_{qp}-J^{u}_{pq}=\frac{e_{0}}{h}|t_{qp}|^{2}(\mu_{p}-0)-\frac{e_{0}}{h}|t_{pq}|^{2}(\mu_{q}-0)$
which is a charecteristic current response for the system due to $\mu_{p}$ and $\mu_{q}$ alone with other probes removed. It is to be noted that $|t_{qp}|^{2}(\mu_{p}-0)$ is a notation for $|t_{qp}|^{2}\Delta\mu_{p}$ and that $|t_{qp}|^{2}(\mu_{p}-0)\neq|t_{qp}|^{2}(\mu_{p})-|t_{qp}|^{2}(0)$. Never the less the zero chemical potentials will not contribute because we expect that for $\mu_{p}=\mu_{q}\neq 0$ we should get $J_{qp}=0$ as a property of the system independent of measurements and hence, $t_{pq}=t_{qp}\neq 0$, which is also what we would get from quantum mechanics of a single particle. For $\mu_{p}=\mu_{q}$ there will be no current in the system even if no measurements are made. Thus
$J_{qp}=\frac{e_{0}}{h}(|t_{qp}|^{2}\mu_{p}-|t_{pq}|^{2}\mu_{q})$ (29)
Comparing with Eq. 14 one has to notice the following differences. In case of Eq. 14, without any loss of generality, we had assumed $(\mu_{1}-\mu_{2})$ is positive, implying current is flowing from lead 1 to lead 2, and that below $\mu_{2}$ there is no current. In the present case with many contacts, $(\mu_{p}-\mu_{q})$ can be positive as well as negative. In Eq. 29, $J_{qp}$ can be positive as well as negative and is a property of the system, that generally may not be measured when there will be many particles and many leads involved. In reality, with only two leads, below $\mu_{2}$, the current form 1 to 2 equals the current from 2 to 1 and this is expected from the principle of detailed balance, and this principle can safely be assumed in Eq. 14. And so for Eq. 29, we are considering the currents all the way down to zero incident energy, that is we are not assuming any detailed balance between two particular reservoirs. Only expectation is that multiple reservoirs will compete to ensure some kind of detailed balance that will come out of our analysis and tested experimentally. Indeed we will see that we will get new kinds of reciprocity relations when there are many leads, that are different from Onsager reciprocity relation that was derived from the principle of detailed balance in the thermodynamic limit. Onsager had shown that for resistivity, $\rho_{xx}(B)=\rho_{xx}(-B)$ and so one would expect that resistance will remain the same on reversing the magnetic field. But signature of this was not found for multi lead mesoscopic systems and the broad answer has to do with the fact that for mesoscopic systems we cannot define a resistivity, as stated in the introduction. For small differences like $(\mu_{p}-\mu_{q})$, we can proceed with linear response. That is we are not considering terms in $(\mu_{p}-0)^{2}$, etc, and higher powers. For only two leads, since the measured resistance is the sample resistance plus the contact resistance, we always get $t_{pq}(B)=t_{pq}(-B)$ as expected on physical grounds, because reversing the magnetic field is just looking at the sample from above or bottom, and the two probe measured conductancd as given by Eq. 14 always exhibit this, giving an impression that there may be a $\rho_{xx}(B)=\rho_{xx}(-B)$ but multiprobes will rule that out. In general $t_{pq}(B)\neq t_{pq}(-B)$ in presense of more leads than $p$ and $q$. For example, in presence of magnetic field $B$ and multiple leads, it is found
$t_{pq}(+B)=t_{qp}(-B)$ (30)
meaning transmission amplitude from $p$ to $q$ is the same as the transmission amplitude from $q$ to $p$ with the direction of magnetic field reversed. Time reversibility of quantum mechanics and duality between wave-function and its conjugate is necessary to explain this [1]. Besides we will see that in sum total (say we sum over q) even for a fixed magnetic field (not reversed), below a certain chemical potential, there will be no current to or from $p$, justifying assuming no current below $\mu_{2}$ for the two terminal case.
Thus Eq. 29 will be valid even in presence of many leads (more than two leads) and if we can write the electrostatic potential in the $p$th reservoir as
$U_{p}=\frac{\mu_{p}}{e_{0}}$ (31)
then
$J_{qp}=\frac{e_{0}^{2}}{h}(|t_{qp}|^{2}U_{p}-|t_{pq}|^{2}U_{q})$
$J_{p}=\sum_{q}J_{qp}=\sum_{q}(G_{qp}U_{p}-G_{pq}U_{q})$ (32)
where
$G_{pq}=\frac{e^{2}_{0}}{h}|t_{pq}|^{2}$ (33)
If we make $U_{q}=U_{p}$ for all $q$ in Eq. 32, then we must get $J_{p}=0$ which gives the following sum rule (which will be explicitely shown below as we discuss three probe conductance).
$0=\sum_{q}(G_{qp}U_{p}-G_{pq}U_{p})$
$\sum_{q}G_{qp}=\sum_{q}G_{pq}$ (34)
This is true for all $p$ and $U_{p}$ and hence is an identity. Eqs. 30 and 34 taken together can be considered new reciprocity relation. That is
$\sum_{q}G_{qp}(+B)=\sum_{q}G_{pq}(+B)$
$G_{qp}(+B)=G_{pq}(-B)$
This is never found to be violated in an experiment. Substituting Eq. 34 into Eq. 32 we get
$J_{p}=\sum_{q}G_{pq}(U_{p}-U_{q})$ (35)
Now if the $p$th lead needs to be used as a voltage probe then for $U_{p}\neq U_{q}$, it should not draw any current or $J_{p}$=0. This condition yields a solution to Eq. 35 giving the following $U_{p}$ which is what would be the voltage at the point where the $p$th lead attaches to the sample.
$U_{p}=\frac{\sum_{q\neq p}G_{pq}U_{q}}{\sum_{q\neq p}G_{pq}}$ (36)
This is true for arbitrary number of $p$ and $q$ and so can be applied to a 3 probe set up or a 4 probe set up, but does not apply to the two probe case.
Let us talk of a three probe conductance instead of a four probe conductance because that is very informative. This means $p$ and $q$ take values from the set $\{1,2,3\}$. Which means lead 1 is from the source reservoir and lead 3 is going into the sink reservoir. That also means $U_{1}$ and $U_{3}$ are applied by an experimentalist while he is trying to measure $U_{2}$ which gives the voltage of the floating point to which lead 2 attaches to the sample. In Eq. 35 let us put $p=1$ and let there be three leads only, which gives.
$J_{1}=G_{11}(U_{1}-U_{1})+G_{12}(U_{1}-U_{2})+G_{13}(U_{1}-U_{3})$
Or
$J_{1}=(G_{12}+G_{13})U_{1}-G_{12}U_{2}-G_{13}U_{3}$
Similarly
$J_{2}=-G_{21}U_{1}+(G_{21}+G_{23})U_{2}-G_{23}U_{3}$
Similarly
$J_{3}=-G_{31}U_{1}-G_{32}U_{2}+(G_{31}+G_{32})U_{3}$
Therefore
$\begin{bmatrix}J_{1}\\ J_{2}\\ J_{3}\end{bmatrix}=\begin{bmatrix}G_{12}+G_{13}&-G_{12}&-G_{13}\\ -G_{21}&G_{21}+G_{23}&-G_{23}\\ -G_{31}&-G_{32}&G_{31}+G_{32}\end{bmatrix}\begin{bmatrix}U_{1}\\ U_{2}\\ U_{3}\end{bmatrix}$ (37)
To cast this in the form of a matrix equation, we define
$G_{12}+G_{13}=-G_{11}$
$G_{21}+G_{23}=-G_{22}$
$G_{31}+G_{32}=-G_{33}$
Hence
$\begin{bmatrix}J_{1}\\ J_{2}\\ J_{3}\end{bmatrix}=\begin{bmatrix}-G_{11}&-G_{12}&-G_{13}\\ -G_{21}&-G_{22}&-G_{23}\\ -G_{31}&-G_{32}&-G_{33}\end{bmatrix}\begin{bmatrix}U_{1}\\ U_{2}\\ U_{3}\end{bmatrix}$ (38)
which gives
$J_{1}=-G_{11}U_{1}-G_{12}U_{2}-G_{13}U_{3}$ (39)
Physically we may expect that if all voltages are the same then we will get no currents. That is if we set $U_{1}=U_{2}=U_{3}\neq 0$ then $J_{1}=J_{2}=J_{3}=0$ must happen. That implies, from Eq. 38
$0=-G_{11}U_{1}-G_{12}U_{1}-G_{13}U_{1}$ (40)
Hence
$G_{11}+G_{12}+G_{13}=0\;\;\mbox{and}\;\;-G_{11}=G_{12}+G_{13}$ (41)
$G_{21}+G_{22}+G_{23}=0\;\;\mbox{and}\;\;-G_{22}=G_{21}+G_{23}$ (42)
$G_{31}+G_{32}+G_{33}=0\;\;\mbox{and}\;\;-G_{33}=G_{31}+G_{32}$ (43)
which justifies the definition of $G_{11}$ and also partly explains Eq. 34. Next we expect Kirchoff’s law to hold, which means $J_{1}+J_{2}+J_{3}=0$ for arbitrary $U_{i}$s.
$J_{1}+J_{2}+J_{3}=0$ (44)
So from Eq. 38 we get
$G_{11}U_{1}-G_{12}U_{2}-G_{13}U_{3}-G_{21}U_{1}-G_{22}U_{2}-G_{23}U_{3}-G_{31}U_{1}-G_{32}U_{2}-G_{33}U_{3}=0$ (45)
$(G_{11}+G_{21}+G_{31})U_{1}-(G_{12}+G_{22}+G_{32})U_{2}-(G_{13}+G_{23}+G_{33})U_{3}=0$ (46)
Since this is true for any arbitrary $U_{i}$s
$G_{11}+G_{21}+G_{31}=0$ (47)
$G_{12}+G_{22}+G_{32}=0$ (48)
$G_{13}+G_{23}+G_{33}=0$ (49)
This justifies the rest of Eq. 34. Note that the $G$s are determined by the system. The voltages and currents are determined by the $G$s. Some of the voltages may be applied and the others are response determined by the response functions $G$s. Now we have descided for the system in Fig. 2, which is a current lead and which is a voltage lead. Lead 2 is the voltage lead which will measure the voltage of the floating point to which it is attached and $U_{2}$ will be adjusted such that $J_{2}$ becomes zero and this adjusted value will give the voltage of the floating point to which lead 2 is attached. While fixed $U_{1}$ and $U_{3}$ is applied to provide a bias that drives a current from 1 to 3, the basic property of a voltage probe is that it should not draw any current. One can choose to set $U_{3}=0$ without any loss of generality and write from Eq. 36.
$U_{2}=\frac{G_{21}}{G_{21}+G_{23}}U_{1}$ (50)
From Eq. 37 we can write for $U_{3}=0$
$J_{2}=-G_{21}U_{1}+(G_{21}+G_{23})U_{2}$ (51)
$J_{3}=-G_{31}U_{1}-G_{32}U_{2}$ (52)
We need not consider $J_{1}$ because by knowing $J_{2}$ and $J_{3}$, we can know $J_{1}$ from Eq. 44. Since the voltage probe does not draw any current, $J_{2}=0$ and so Eq. aaaa gives back Eq. aaaa. From Eq. aaaa for $J_{3}=J_{1}$ we get
$\frac{J_{1}}{U_{1}}=-G_{31}-G_{32}\frac{G_{21}}{G_{21}+G_{23}}$ (53)
This is always arbitary to a sign depending on convention that current entering the system is positive and current leaving the system is negative or vice versa. In fact Eq. 50 should be written as $J_{1}+J_{2}-J_{3}=0$. Or we can also express it as
$\frac{J_{1}}{U_{2}}=\frac{G_{21}+G_{23}}{G_{21}}\bigg(-G_{31}-G_{32}\frac{G_{21}}{G_{21}+G_{23}}\bigg)$ (54)
Since Eqs. 53 and 54 has been experimentally verified time and again, in presence or absence of magnetic fields, one can say that Pauli exclusion principle associated with LPDOS has also been verified experimentally.
2.2 Demonstration of $\rho_{lpd}$ from $G_{\alpha\beta}$s
While Eqs. 53 and 54 is measured by making $J_{2}=0$ by adjusting $\mu_{2}$, one can also earth the STM tip as shown in Fig. 3 and measure the coherent current from reservoir 1 to reservoir 2. This coherent current can demonstrate $\rho_{lpd}$ as will be shown in this subsection. This coherent current can be again measured by measuring scattering probabilities as is evident from Eq. 14. In axiomatic quantum mechanics, any local state will be disturbed by measurements in an unpredictable manner and the theory of quantum mechanics rests on the fact that there can only be a random outcome. There is no theory for this measurement related randomness in quantum mechanics, as opposed to the dynamics prior to measurements for which probabilities can be calculated. Once again Buttikker and his coworkers [8] could demonstrate a connection between measured currents and LPDOS in the semi-classical regime. The real issue is in the quantum regime and that remained puzzling as LPDOS could become negative and they could not make sense of it. We had shown that negative LPDOS is consistent with negative signal propagation time or time travel [2]. In the previous subsection we also showed that LPDOS can answer the ananswered questions related to why Landauer-Buttiker formalism works so well. In this sub-section, we will demonstrate that the relation between LPDOS and measured scattering probability, as derived by Buttikker and co-workers make perfect sense even when we get negative LPDOS. It will be thus clear that the possibility of a deterministic local state with a LPDOS also solve the quantum measurement problem as will be demonstrated in this sub-section. Just as in principle, one can solve Newtons equation for $10^{23}$ particles to derive classical statistical mechanics, one can in principle calculate LPDOS for a mesoscopic sample with $10^{23}$ leads attached, to deterministically predict every single electron scattering outcome. As to why the uncertainty principle will not affect this deterministic outcome is discussed in [2] (see Eqs. 3.56 and 3.57 therein with the discussion below it), but the intuitive answer is that local states do not appear in axiomatic quantum mechanics but the axioms do not rule out its possibility by Goedel’s incompleteness theorem. This applies to any other theorem that can be proved within the axiomatic framework of quantum mechanics, like for example the Bell’s theorems.
For the purpose of this demonstration we consider the three probe (three leads attached to a mesoscopic sample) set up as described below with respect to Fig. 2, but the logic can be extended to arbitrary number of leads. The mesoscopic sample is the shaded region which is typically made of semi-conductor or metal. The leads $\gamma$ or $\alpha$ are current leads that may be made up of the same material. Lead $\beta$ is an STM tip that can attach to different points in the sample and is such that we can vary its proximity to the sample. The STM tip can be set up in many different ways but we will focus on the particular case considered by Buttiker and his co-workers [8]. It will be clear as to why they could only understand the semi-classical regime of that set up. That is the set up when the STM tip is weakly coupled to the sample at a point r through a tunnelling barrier.
Figure 2: Fig. 2. The basic set up for a three probe Landauer conductance where there are only two fixed leads indexed $\gamma$ and $\alpha$ apart from the STM tip $\beta$.
Figure 3: Fig. 3. The sample is the three prong potential shown by the solid lines and the entire system consist of the sample connected to three reservoirs via three leads. Different regions of the system is marked by Roman numbers, I, IV and VII being the leads, shown by dashed lines. Lead 2 is similar to the lead $\beta$ in Fig. 2 but earthed and the chemical potential of lead 3 is also set to zero. In that case, conventionally, a small positive chemical potential at lead 1 will inject a current while the two other leads carry current away from the sample. Lead 2 connects to the sample through a tunneling barrier shown as region VI. Lengths of different regions is shown as $l_{2}$, $l_{3}$ etc. This system is a 1D version of the system in Fig. 2 where lead $\gamma$ is renamed as lead 1, etc, and as a result $t_{31}\equiv s_{31}$, $t_{21}\equiv s_{21}$, and $r_{11}\equiv s_{11}$.
Let chemical potential of left reservoir connected to lead $\gamma$ be $\mu_{\gamma}>0$, chemical potential of right reservoir connected to lead $\alpha$ is $\mu_{\alpha}=0$, and the other end of the the lead $\beta$ is earthed. Essentially, $\mu_{\gamma}$ and $\mu_{\alpha}$ are fixed for all incident energies (or other parameters), but $\mu_{\beta}$ can floatingly adjust to zero. For incident energy $E$ such that $0=\mu_{\alpha}<E<\mu_{\gamma}$, only lead $\gamma$ injects a current according to our classical notions. For a mesoscopic sample, according to Landauer-Buttiker formalism, this defines a quantum mechanical scattering problem. Such a set up can be obtained in the laboratory and we want to address the coherent current flowing from $\gamma$ to $\alpha$. One can see [3] that in such a situation
$|s_{\alpha\gamma}^{\prime}|^{2}-|s_{\alpha\gamma}|^{2}=-4\pi^{2}\mid{t_{\beta}}\mid^{2}\nu_{\beta}\rho_{lpd}(E,\alpha,\textbf{r},\gamma)$ (55)
where
$\rho_{lpd}(E,\alpha,\textbf{r},\gamma)=-\frac{1}{2\pi}\;|s^{\prime}_{\alpha\gamma}|^{2}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$ (56)
To make a connection with Eq. 1, it is suitable to define
$\rho_{lpd}^{r}=\rho_{lpd}(E,1,\textbf{r},1)\;\;\mbox{and}\;\;\rho_{lpd}^{t}=\rho_{lpd}(E,3,\textbf{r},1)$
which are local partial density of states (LPDOS) for this system, where $U(\textbf{r})$ is the electrostatic potential at the point r,
$\theta_{s^{\prime}_{\alpha\gamma}}=ArcTan\bigg[\frac{Im[s^{\prime}_{\alpha\gamma}]}{Re[s^{\prime}_{\alpha\gamma}]}\bigg]$ (57)
and $\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$ is a functional derivative. The parameter $t_{\beta}$ controles coupling of states of the sample with the states at the STM tip for which $\nu_{\beta}$ is the relevant density of states (DOS). For a commercially available STM tip $t_{\beta}$ and $\nu_{\beta}$ are just parameters, either specified by the manufacturer or to be ingeniously approximated by the user. Thus, $s_{\alpha\gamma}^{\prime}$ is the scattering matrix element for scattering from $\gamma$ to $\alpha$ when the STM tip is drawing a current given by the RHS of Eq. 55. When the STM tip is removed by $t_{\beta}\rightarrow 0$ then this scattering amplitude will be $s^{\prime}_{\alpha\gamma}=s_{\alpha\gamma}$. Coherent currents are given by $\frac{e_{0}}{h}|s^{\prime}_{\alpha\gamma}|^{2}$ and $\frac{e_{0}}{h}|s_{\alpha\gamma}|^{2}$ in the two situations.
In earlier works, Eq. 55 was studied in the semi-classical regime wherein $\rho_{lpd}$ is positive definite, which limits its value and applicability in the quantum regime. The quantum regime remained unclear specially because in the quantum regime $\rho_{lpd}$, as caclculated from Eq. 56, can become negative. We will show that Eq. 55 remains valid even when $\rho_{lpd}$ becomes negative in a quantum regime. For this it is extremely important to apply wave-mechanics or von Neumann quantum mechanics for the region between the two reservoirs, that is the mesoscopic system. The key idea needed to show this is to realize the fact that in the quantum regime of 1D, 2D and 3D, $\rho_{lpd}$ cannot really be negative although sometimes it misleads one to think so. Real negativity of $\rho_{lpd}$ is a speciality of the mesoscopic regime when we encounter Fano resonances [2]. This is the primary reason why earlier works could not understand the quantum regime and to be demonstrated below.
Intuitively speaking, $\rho_{lpd}$ is positive and the LHS of Eq. 55 is negative accounting for loss of coherent electrons to the earthed lead $\beta$. Therefore, any change of sign in $\rho_{lpd}$ will counter-intuitively result in a gain of coherent current due to the intervention of the earthed STM tip on a quantum state. That can be regarded as a supporting evidence for time time travel (like a quantum state getting younger with respect to coordinate time) because any intervention on a quantum state at an intermediate point, by an STM tip is supposed to only decohere the state, with measurement outcomes being random. We will clarify as to why and when such counterintuitive effects can be seen in real systems. Since LHS of Eq. 55 can be defined in quantum mechanics and can also be measured in the laboratory and so if the equality in Eq. 55 can be established then one can affirm that RHS is an objective reality that can be found in nature, although we cannot make sense of it within the axiomatic framework of quantum mechanics. The object $\rho_{lpd}$ cannot be defined within the axiomatic framework of quantum mechanics because it is refering to only those electrons that are coming from $\gamma$ and going to $\alpha$. The wave-function of an electron at a point r can only be sensitive to either $\gamma$ or $\alpha$. But the object can be defined for a mesoscopic system by appealing to the idea of a physical clock.
If we start from a closed system govorned by Schrodinger equation, we cannot get an object that is $\gamma$ and $\alpha$ dependent because in that case the idea of a physical clock do not work as they inherently need asymptotic free states [2]. In a mesoscopic set up these asymptotic states can not be in a linear superposition. The classical reservoirs destroy the linear superposition of states in the leads while linear superposition exists inside the sample (shaded region) [2]. Events in the reservoir are all classical events and quantum event can only be defined in between two classical events in the reservoirs. This will be discussed again with respect to the 1D picture given in Fig. 3. Once we know $\rho_{lpd}$ we can use it to find other relevant quantities like injectance that we find in axiomatic framework of quantum mechanics and so we are not loosing anything but only gaining some extra objects by this assertion. At any point we can go to the closed system by making the leads $\gamma$ and $\alpha$ very weakly connected to the system and recover all that we get by solving the Schrodinger equation for the closed system. Besides it may not be always necessary to pinch off the leads physically as von Neumann quantum mechanics give us a method to infer about the closed system from the scattering states as has been worked out in chapter 4 of Ref [2].
In this paragraph we discuss that $\rho_{lpd}$ as a local object solves the problem of collapse of wave-function by making it redundant, with measurements defined deterministically in the classical world of the reservoirs, while the time evolution of observables in the region between the reservoirs is determined by quantum mechanical scattering. At a point r in the shaded region, there will be electrons in a linear superposition of states. One may say that all these electrons came from lead $\gamma$ (assume for the time being that $\rho_{lpd}$ is positive definite) but some may be going to lead $\alpha$ (we call them dancing cats), some may be going to lead $\beta$ (we call them sleeping cats) and some may be getting reflected back to lead $\gamma$ (we call them beaten cats). At the point r, we have a linear superposition of dancing, sleeping and beaten cats. But we can now say something about only the dancing cats. They spend a time $\frac{h}{2\pi}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$, at the point r. At zero temperature there are $|s^{\prime}_{\alpha\gamma}|^{2}$ of these dancing cats at that point and we can average over only the dancing cats to define a local partial density of states given by $\rho_{lpd}(E,\alpha,\textbf{r},\gamma)=-\frac{1}{2\pi}\;|s^{\prime}_{\alpha\gamma}|^{2}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$. That means we can selectively average over the dancing cats at the point r and determine physical effects due to the dancing cats alone. In formal quantum mechanics we cannot do this selective averaging for the dancing cats alone that are in a linear superposition with the sleeping and beaten cats. Note that $|s^{\prime}_{\alpha\gamma}|^{2}$ and $\theta_{s^{\prime}_{\alpha\gamma}}$ are measured at the junction of the sink reservoir which is a classical system. Probability is well defined for classical systems and measured $|s^{\prime}_{\alpha\gamma}|^{2}$ is a probability wherein measurements do not care how the probability is defined theoretically. Similarly, according to the notion of physical clock, $\theta_{s^{\prime}_{\alpha\gamma}}$ is defined as angular displacement of a precessing dipole. A perfectly logical connection can be made between quantum mechanical scattering phase shift $\theta_{s^{\prime}_{\alpha\gamma}}$ and well defined angular displacement of a precessing dipole [2]. Thus the measurables are not random outcomes induced by collapse of wavefunction by the intevening STM tip at r, but can be calculated precisely in the Landauer-Buttiker formalism wherein the punchline is that a quantum event can only be defined between two classical events in the reservoirs that perfectly help define a local time and a local state at the point r. This will be demonstrated with the help of Eq. 55 wherein the LHS has the measurables and the RHS has the local object. We know that von Neumann quantum mechanics has one mathematical framework for scattering states and bound states. For any scattering phenomenon the LPDOS can be defined and bound states can be seen as a special case of the scattering states. So we believe that this solves the measurement problem in general and we have to see how the Landauer-Buttiker formalism can be extended to other practical cases. At various places in this work, we have sketched out this roadmap as to what is a general mesoscopic system and how any other system can be seen as a modification of it. That is why we have discussed von Neumann formalism, Diracs approach, Heisenberg’s approach, S. Datta’s approach, reservoirs being in the domain of relativity, etc.
Thus we intend to justify Eq. 55 for the quantum regime, specially the situation when $\rho_{lpd}$ becomes negative. Ref. [8] gives a derivation of Eq. 55, we will give a physical topological arguement here which will make it clear why and when we can generally expect the equality in Eq. 55 inspite of $\rho_{lpd}$ being negative. That will be followed by a numerical demonstration. Note that there are two DOS that feature in Eq. 55, that are $\nu_{\beta}$ and an r dependent DOS that is $\rho_{lpd}$. There is a factor $\frac{1}{2\pi}$ for each of them that cancels the factor $4\pi^{2}$. This is the $\frac{1}{2\pi}$ factor that appears in Eq. 56. There can be a very simple interpretation for this factor as follows. LHS of Eq. 56 is a DOS that can accommodate electrons that are countable but $2\pi\rho_{lpd}$ is uncountable, the RHS of Eq. 56 being in terms of $\theta_{s_{\alpha\gamma}}(E)$ corresponding to a continuous rotation. This also make it essential that if only one half of a Riemann surface is involved then rotaion angle can only be maximum $\pi$ instead of $2\pi$. So winding number in Eq. 56 as a conserved quantity in the two cases will differ by a factor of 2. On the other hand, scattering probabilities $|s^{\prime}_{\alpha\gamma}|^{2}$ and $|s_{\alpha\gamma}|^{2}$ are just numbers, are well defined in quantum mechanics, and when multiplied by a factor $\frac{e_{0}}{h}$ they give the measured coherent current from $\gamma$ to $\alpha$.
Let us consider the system shown in Fig. 3. First of all it is a system for which exact quantum mechanical calculations can be made and it is a system that exhibit pronounced Fano resonances. The system in Fig. 2 will generally show such resonances was shown in earlier works. In this sense the system in Fig. 2 at Fano resonances with the modification that the upper lead is weakly coupled to earth via a strong potential can be represented by the 1D system in Fig. 3. Besides, since the logic, to be presented below, is based on topology of the Argand diagram (AD) at a resonance and the complex plane, there are only two kinds of behavior, one being typically like that of Breit-Wigner resonance that involve the full Riemann surface, and the other being typically like that of Fano resonance that involve one half of the Riemann surface. Also since it is a question of whether LPDOS is present at all in a quantum system and more importantly whether its negativity is allowed in quantum mechanics, then demonstrations in 1D quantum mechanics should suffice. Let us drop the factor $|t|^{2}\nu_{\beta}$ from Eq. 55 because such factors will be only needed if we are using a commercially available STM tip, where this factor characterizes the STM tip. This require us to drop a $2\pi$ (or $\pi$) factor in Eq. 55 for reasons mentioned in the previous paragraph. Which means for a first principle calculation for the system in Fig. 3 we expect (if the equality in Eq. 55 is correct) that
$|s_{\alpha\gamma}^{\prime}|^{2}-|s_{\alpha\gamma}|^{2}=-2\pi\rho_{lpd}(E,\alpha,\textbf{r},\gamma)$ (58)
Now to make it correspond to the notations in Fig. 3
$|t_{31}^{\prime}|^{2}-|t_{31}|^{2}=-2\pi\rho_{lpd}(E,3,\textbf{r},1)$ (59)
Figure 4: Fig. 4. In this figure we plot the AD for $t_{31}$ of the system shown in Fig. 3 for the coupling potential $U_{1}$ varying in a range that give three sub-loops, all within one Riemann surface. The starting point is marked by a small square block corresponds to $U_{1}l=-10000$. The end point is unmarked and corresponds to a value $U_{1}l=-10$. All the sub-loops smoothly come back to a point marked by a cross. Other parameters are mentioned inside the figure.
In Fig. 4 we plot the Argand diagram (AD) for $t_{31}$ for the system shown in Fig. 3 by varying the tunneling potential $U_{1}$ which connects the three prong potential with lead 2 that is earthed. As $U_{1}$ is varied from $e_{0}U_{1}l=-10$ to $e_{0}U_{1}l=-1000$ we plot $Im[t_{31}]$ versus $Re[t_{31}]$, which is known as AD. The AD diagram make smooth sub-loops within one Riemann surface and each sub-loop is due to a Fano resonance. Any physical quantity that depend on $t_{31}$ will go through a cycle over one particular sub-loop. For one closed sub-loop generated by monotonously varying a parameter, say $U_{1}(\textbf{r})$, the local potential at r.
$\oint_{c}\Delta\theta_{t_{31}}=\oint_{c}\frac{\delta\theta_{t_{31}}}{\delta U_{1}({\bf{r}})}\;\Delta U_{1}({\bf{r}})=0\;\;\mbox{where}\;\;\theta_{t_{31}}=Arctan\frac{Im[t_{31}]}{Re[t_{31}]}$
Because the AD was generated by varying $U_{1}(\textbf{r})$, the integration measure is $\Delta U_{1}(\textbf{r})$ and the integration for a closed loop will be zero. Or
$\oint_{c}{\frac{1}{2\pi}|t_{31}|^{2}\;\frac{\delta\theta_{t_{31}}}{\delta U_{1}({\bf{r}})}}\;\Delta U_{1}({\bf{r}})=-\oint_{c}{\rho_{lpd}(E,3,\textbf{r},1)}\;\Delta U_{1}({\bf{r}})=0$ (60)
Similarly over the same sub-loop,
$\oint_{c}\frac{\delta|t_{31}|^{2}}{\delta U_{1}(\textbf{r})}\Delta U_{1}(\textbf{r})=\oint_{c}{\Delta|t_{31}|^{2}}=\oint_{c}{\Delta|Re[t_{31}]+Im[t_{31}]|^{2}}=0$ (61)
Therefore from Eqs. 60 and 61 we get the equality
$\oint_{c}{\Delta|t_{31}|^{2}}=\oint_{c}{\frac{1}{2\pi}|t_{31}|^{2}\;\frac{\delta\theta_{t_{31}}}{\delta U_{1}({\bf{r}})}}\;\Delta U_{1}({\bf{r}})=-\oint_{c}{\rho_{lpd}(E,3,\textbf{r},1)}\;\Delta U_{1}({\bf{r}})=0$ (62)
purely as a consequence of the topology of a complex plane $(Im[t_{31}],Re[t_{31}])$. In this complex plane $(Im[t_{31}],Re[t_{31}])$, the point (0,0) is a singular point where the phase $\theta_{t_{31}}=Arctan\frac{Im[t_{31}]}{Re[t_{31}]}$ is undefined and is well known for classical waves [9]. If the Argand diagram trajectory would have enclosed the singularity, then the RHS of Eq. 60 would have been $2\pi$ instead of zero but the RHS of Eq. 61 would be still zero, in which case Eq. 62 will not be valid. We get Eq. 62 from Eqs. 60 and 61 because due to Fano resonances, one gets loops in the AD within one Riemann surface. Whatever be the terms in the Hamiltonian or the equation of motion, it is the topology of the complex plane that determines the outcome of the integration in Eqs. 60 and 61. For the kind of Argand diagram in Fig. 4, that are smoothly closed within one Riemann surface, the equality in Eq. 62 is for the integrals. We will address the value of the integrands latter but the integrands will go through a positive-negative cycle over a loop simply because of the fact that the integration over a loop will be zero and there is no surprize if $\rho_{lpd}$, which is the integrand in Eq. 60, becoming negative. However, the question remains if it is physical and if it can be found in nature. That can be verified by also considering the integrand in the LHS of Eq. 62.
Therefore, it is obvious from Fig. 4 and Eq. 62, that $\Delta|t_{31}|^{2}$ as well as $\rho_{lpd}$ will go through a positive-negative cycle over a closed loop as a consequence of the topology of the relevant complex plane, only if the AD closes within one side of the Riemann surface, as in the case of Fig 4, the AD is restricted to the first and the fourth quadrant. Given the fact that cycles in $\Delta|t_{31}|$ is physical, the cycles of $\rho_{lpd}$ is also physical as they are just different expressions of the same loop made by the AD of $t_{31}$. If we want to compare the integrands in Eq. 62, then we need an extra factor of $2\pi$ as in Eq. 58 and 59 as derived in [8]. Thus
${\Delta|t_{31}|^{2}}\approx 2\pi{\frac{1}{2\pi}|t_{31}|^{2}\;\frac{\delta\theta_{t_{31}}}{\delta U_{1}({\bf{r}})}}\;\Delta U_{1}({\bf{r}})$ (63)
Viewing a derivative as an effect of an infinitesimal change in $U_{1}$ we get
$|t^{\prime}_{31}|^{2}-|t_{31}|^{2}\approx|t^{\prime}_{31}|^{2}(\theta_{t^{\prime}_{31}}-\theta_{t_{31}})$ (64)
where primed quantities and unprimed quantities are calculated for an infinitesimal difference of $U_{1}$. If we want to compare integrands then we can only write an approximate equality because we are comparing two different objects over one sub-loop. There will always be in the least a phase difference between the two quantities. But if $\rho_{lpd}$ determines the measured quantity $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ then the two sides of Eq. 64 should agree in amplitude and frequency. The agreement in frequency will definitely be there because that is determined by the same AD.
In Fig. 5 we plot LHS and RHS of Eq. 64 where the primed values and unprimed values are again for small differences in $U_{1}$ as the incident wave vector $k$ is varied and that is the horizontal axis. Initially the two curves are a little different but as $k$ increases, effects of dispersion is minimized, and magnitudes of the two curves are similar with a phase difference between them. The oscillations of both the curves are due to the smooth cyclic nature of the AD in Fig 4. In other words, the underlying principle behind the oscillation of both curves is the same. The two curves oscillate as a result of quantum interference. But the symmetric oscillation between positive and negative values and the similarity between the two curves can only be noted if there are well defined Fano resonances and signify the fact that the RHS of Eqs. 58 and 59 are physical. For Breit-Wigner resonances when the AD starts going around the singular point there will be deviations. Since Buttiker et al did not consider Fano resonances, the negativity of $\rho_{lpd}$ is unphysical and they did not understand Eqs. 58 and 59 in the quantum regime.
Figure 5: Fig. 5. In this figure we are plotting the LHS and RHS of Eq. 64 to show that they both can oscillate between positive and negative values. The primed and unprimed values are for small differences in $U_{1}$ at a $k$ value that we vary continuously. The sign change and magnitude of both the curves originate from the smooth cyclic AD of Fig. 4.
3 Conclusion
AD of a scattering matrix element like $t_{31}$ or $s_{\alpha\gamma}$ can show two kinds of topological behavior. In one case the AD closes on one side of the complex plane $(Im[t_{31}],Re[t_{31}])$ without enclosing the singularity (0,0). In that case both $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ will oscillate with positive-negative values and will be equal in magnitude. Thus the equality derived in Ref. [8] make sense and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ is determined by $\rho_{lpd}$. However, if the AD of $t_{31}$ does enclose the singularity in the plane $(Im[t_{31}],Re[t_{31}])$ then the equality of the integrations in Eq. 62 itself breaks down because of topological reasons. Then the integrands will still be comparable in magnitudes as derived in Ref. [8] but they will be qualitatively different, not resembling each other. The LHS of Eq. 62 not being a winding number will always be zero while the RHS being a winding number will be $2\pi$. Never the less, $\rho_{lpd}$ still determines $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ as they are both expressions of the same AD, however $\rho_{lpd}$ will oscillate on the positive side only while $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ will oscillate with positive and negative values. One can experimentally determine this AD [10] and by carefully examining the AD one can infer the behavior of $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ and relate them. However, it is easy to tune a mesoscopic system into a regime with only Fano resonances and thereby obtain an equality between $\rho_{lpd}$ and $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ for any system. Thus in principle a measurable quantity $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ is determined by a number of local partial states $\rho_{lpd}(\textbf{r})$ inspite of all possible linear superposition of states at the point r is demonstrated and hence solves the quantum measurement problem in the extreme quantum regime as well. Since $\rho_{lpd}(\textbf{r})$ is thereby real, and is related to a local time that is fully consistent with relativity as well as quantum mechanics, hence unifies the two. Prof. Roger Penrose has been repeatedly stating in his public talks that since outcome of a quantum measurement within the axiomatic framework of quantum mechanics is random it will not be possible to unify relativity and quantum mechanics unless one can show that quantum measurements are deterministic. That precisely come out in our work. The $\rho_{lpd}$ cannot be defined in the axiomatic framework of quantum mechanics, but can deterministically give the outcome of a quantum measurement, and naturally solve the problem of unification.